i have two questions that are very similiar..
$8,000 is compounded annually for 3 years at a rate of 6%. How much is in the account at the end of the term?
$4,000 compounded semiannually for 5 years at a rate of 4%. How much is in the account at the end of the term?

thank you for your time!

Answer by stanbon(75887)

Amount after t years when n is the # of times compounded each year and rate=r
---------------------
A(t)=P(1+(r/n))^(nt)
---------------------------
$8,000 is compounded annually for 3 years at a rate of 6%. How much is in the account at the end of the term?
A(3) = 8000(1+ (0.06/1)^(1*3)
= 8000(1.06)^3
= 8000*1.1910..
A(3) = $9528.13
------------------

4,000 compounded semiannually for 5 years at a rate of 4%. How much is in the account at the end of the term?
A(5) = 4000(1+(0.04/2))^(2*5)
A(5) = $4875.98
========================
Cheers,
Stan H.